OpenStudy (ehall):

A telephone pole 13 1/3 feet tall casts a shadow of 16 feet when a person casts a shadow of 6 feet. How tall is the person? 4 1/3ft 5ft 5 1/3ft

OpenStudy (revlogic):

This is a proportion problem. So let's two proportions equal to each other since we know the ratio between height and shadow are the same. 13.333/16 = x/6 *cross multiply 79.999 = 16x *now solve for x by dividing 16 both sides of the equation 79.999/16 = (16/16)x x = 4.999 or 5 ft

OpenStudy (ehall):

oh that makes sens now thank you so much for helping me and showing me how to do it

OpenStudy (tkhunny):

Never "cross-multiply". That doesn't mean anything. You had this: \(\dfrac{13.333}{16} = \dfrac{x}{6}\) If you "cross-multiply", which doesn't mean anything, you have done two operations: 1) Multiple both sides by 6 and 3) Multiply both sides by 16. This leaves: \(80 = 16x\) The solution, after "cross-multiplying", is obtained by dividing by 16. WAIT!! Didn't we just Multiply by 16? Well, THAT was a waste of time. Just multiply. That's all. You had this: \(\dfrac{13.333}{16} = \dfrac{x}{6}\) Multiply by 6. \(\dfrac{13.333}{16}\cdot 6 = x = 5\) -- Done.

OpenStudy (ehall):

oh that seems way easier then the other way thank you so much

OpenStudy (tkhunny):

Learn the basic operations. "cross-multiply" is NOT one of them.